∫x^2 sqrt(2-x)dx


1
Expert's answer
2021-03-02T00:55:31-0500

Let "u=2-x" , then "-du=dx" and "x=2-u"


Use the above substitution to evaluate the integral as,


"\int x^2 \sqrt{2-x}dx=\int (2-u)^2 \sqrt{u}(-1)du"


"=-\int (4+u^2-4u) \sqrt{u}du"


"=-\int (4\sqrt{u}+u^2\sqrt{u}-4u\sqrt{u})du"


"=-[4({\frac{2}{3}u^\frac{3}{2}})+{\frac{2}{7}u^\frac{7}{2}}-4({\frac{2}{5}u^\frac{5}{2}})]+C"


"=-\frac{8}{3}u^\frac{3}{2}-\frac{2}{7}u^\frac{7}{2}+\frac{8}{5}u^\frac{5}{2}+C"


"=-\frac{8}{3}(2-x)^\frac{3}{2}-\frac{2}{7}(2-x)^\frac{7}{2}+\frac{8}{5}(2-x)^\frac{5}{2}+C"



Therefore, the integral is "\int x^2 \sqrt{2-x}dx=-\frac{8}{3}(2-x)^\frac{3}{2}-\frac{2}{7}(2-x)^\frac{7}{2}+\frac{8}{5}(2-x)^\frac{5}{2}+C"